# on dextral symmetric algebras

### v1

@@ 5. Classification of complex dextral symmetric Leibniz algebras upto dimension $4$
There is no dextral symmetric Leibniz algebra with nilradical isomorphic to $\mu_{4}$ or $\mu_{5}$.
Following the discussions above, we have a complete characterisation of a $4$-dimensional dextral symmetric Leibniz algebra.
@@
A $4$-dimensional solvable Leibniz algebra $\mathscr{L}$ is dextral symmetric if and only if $\mathscr{L}^{(3)}=0.$
@@
We bring this paper to a close with an example of a $7$-dimensional Lie algebra to shows that not every finite-dimensional dextral symmetric Leibniz algebra $\mathscr{L}$ will satisfy $\mathscr{L}^{(3)}=0.$
@@ Acknowledgements
We express our heartfelt gratitude to Dr. Ardeline Mary Buhphang (NEHU, Shillong), Rishabh Goswamy (NEHU, Shillong) and Dr. Ripan Saha (Raiganj University) for their valuable suggestions while preparing this manuscript. The first author would like to thank Prof. Samuel Lopes (University of Porto) for his insightful conversations during CIMPA School-2024, held in Philippines.

### v2

@@ 5. Classification of complex dextral symmetric Leibniz algebras up to dimension 4
There is no $4$-dimensional dextral symmetric solvable Leibniz algebra with nilradical isomorphic to $\mu_{4}$ or $\mu_{5}$.
Following the discussions above, we have a complete characterization of a $4$-dimensional solvable dextral symmetric Leibniz algebra.
@@
A $4$-dimensional Leibniz algebra $\mathscr{L}$ is dextral symmetric if and only if $\mathscr{L}^{(3)}=0.$
@@
We bring this paper to a close with an example of a $7$-dimensional Lie algebra to show that not every finite-dimensional dextral symmetric Leibniz algebra $\mathscr{L}$ will satisfy $\mathscr{L}^{(3)}=0.$
@@ Acknowledgements
We express our heartfelt gratitude to Dr. Ardeline Mary Buhphang (NEHU, Shillong), Rishabh Goswami (NEHU, Shillong) and Dr. Ripan Saha (Raiganj University) for their valuable suggestions while preparing this manuscript. The first author would like to thank Prof. Samuel Lopes (University of Porto) for his insightful conversations during CIMPA School-2024, held in Philippines. We also thank Ivan Kaygorodov’s (Federal University of ABC, Brazil) for identifying an error in our paper and bringing it into our attention.